1. A trolley of mass 3 kg moves with a velocity of 1.5 m s − 1 towards a stationary trolley of mass 2 kg. after the collision, the two trolleys move together. what is the common velocity of the two trolleys?
  2. A pistol weighing 1.4 kg fires out a bullet weighing 1.6 g. if the pistol jerks backwards with a velocity of 0.2 ms-1, what is the velocity of the bullet?
  3. A trolley of mass 400g moving with a velocity of 2.0 ms-1 collides with a trolley B of mass 500g moving in the opposite direction with a velocity of 2.5 ms-1. if trolley B moves with a velocity of 1.5ms-1 in the same direction after the collision, what is the velocity of trolley A after the collision?
  4. An arrow weighing 150g is shot at a piece of wood weighing 450g that is on a slippery surface. when the collision of the arrow moves horizontally with a velocity of 15 ms-1. calculate the common velocity after the collision.
  5. A bullet of mass 20g is fired from a rifle of mass 5kg with a velocity of 200 ms-1. what is the jerk velocity of the rifle?
1
Expert's answer
2021-03-29T09:09:34-0400

1)


"(3)(1.5)=(3+2)v\\v=0.9\frac{m}{s}"

2)


"(1.4)(0.2)=(0.0016)v\\v=175\frac{m}{s}"

3)


"(0.5)(2.5)-(0.4)(2)=(0.5)(1.5)+(0.4)v\\v=-0.75\frac{m}{s}"

4)


"(150)(15)=(150+450)v\\v=3.75\frac{m}{s}"

5)


"(0.02)(200)=(5)v\\v=0.8\frac{m}{s}"